The energy of the lowest allowed level for an electron in a 1-D box is 10.0 eV. Can the electron ever have an energy of 200 eV? Justify.
1
Expert's answer
2019-03-20T11:46:09-0400

The allowed energies for an electron in a 1-D box

"E_n=\frac{n^2\pi^2\hbar^2}{2mL^2}, \quad n=1,2,3,..."

So

"\frac{E_n}{E_k}=\frac{n^2}{k^2}"

"\frac{200\:\rm{eV}}{10.0\:\rm{eV}}=\frac{n^2}{1^2}"

"n=\sqrt{20.0}=4.47"

Since n is non-integer, so the electron would not have an energy of 200 eV.